# PPTA Collaboration

### v1

@@ I Introduction
Pulsars are also known as excellent astronomical linear polarization sources. Their light can be linearly polarized to well-above $10\%$ of total intensity. The PTA programs usually measure the pulsar polarization profiles for calibrating the relevant observations Osłowski et al. (2013). Vast amounts of polarization data thus exist or are expected to be collected in the active PTAs and other pulsar-monitoring programs. This has motivated the suggestion of developing Pulsar Polarization Arrays (PPAs) Liu et al. (2023), to explore astrophysics and fundamental physics with a correlated polarization signal across the pulsars. One representative scientific case is detecting ultralight axion-like dark matter (ALDM).
@@
In this letter, we conduct the first-ever PPA analysis to search for the ultralight ALDM, using the polarization data of 22 MSPs from PPTA Data Release 3 (DR3) Zic et al. (2023); Reardon et al. (2023). We denote this PPA as the Parkes PPA (PPPA) for the convenience of discussions. To accomplish this task, we develop an advanced analysis framework which involves precise construction of pulsar PA residual time series, systematic modeling of their noise contributions, and application of Bayes statistical method with pulsar cross-correlation implemented. Using this framework, we demonstrate that the PPPA provides the most stringent constraints on the ALDM Chern-Simons coupling within the relevant mass range. Notably, we demonstrate for the first time that the cross-correlation among the pulsars plays a crucial role in investigating the nature of the derived limits.
@@ II ALDM induced cosmic birefringence
×$\times\cos[m_{a}(t-\mathbf{v}\cdot\mathbf{x})+\phi_{\mathbf{v}}]\,,$
where $\rho(\mathbf{x})$ is ALDM energy density at position $\mathbf{x}$, $m_{a}$ is ALDM mass, $\mathbf{v}$ denotes velocity in a latticed phase space $\Omega$, $\Delta v$ is lattice spacing, and $f_{\mathbf{x}}(\mathbf{v})$ represents DM velocity distribution at $\mathbf{x}$. The profile of $\rho(\mathbf{x})$ is determined by Schrödinger-Poisson equations. As shown in Schive et al. (2014), the halo self-gravitation leads to the formation of a central solitonic core within galaxies. Since most PPPA MSPs are only $\lesssim\mathcal{O}(10^{3})$ parsec away from the Earth, we simply assume $\rho(\mathbf{x})$ to be uniform, i.e., $\rho(\mathbf{x})=\rho_{0}$ foo . $\alpha_{\mathbf{v}}\in(0,+\infty)$ is an amplitude measure and $\phi_{\mathbf{v}}\in(0,2\pi)$ is a phase parameter. They follow Rayleigh and uniform distributions, respectively Foster et al. (2018), which reflects the stochastic nature of the ALDM field.
@@
where the two terms are dubbed “pulsar” and “Earth” terms. Note, while the SGWB-induced timing residual in the PTA is also determined by so-called “pulsar” and “Earth” terms, unlike the CB effect, it arises from a cancellation between contraction and expansion of the light path as the GWs pass by. Moreover, the CB is distinguished from the Faraday rotation in interstellar medium (ISM) and Earth’s ionosphere as the latter effect is frequency-dependent and relies on the length of light path aligned with the direction of magnetic field.
@@
In last decade, a few attempts have been made to explore temporal variations of pulsar polarization Yan et al. (2011); Wahl et al. (2022). The PA residuals of individual pulsars have been suggested for testing the ALDM theory Liu et al. (2020); Caputo et al. (2019). However, a systematic analysis framework of the PPA, which emphasizes the data correlation across the pulsars, remains absent. To address this, we model the observed PA residuals as:
$\Delta{\rm PA}^{\rm obs}=\Delta{\rm PA}^{a}+\Delta{\rm PA}^{\rm ion}+\Delta{\rm PA}^{w}+\Delta{\rm PA}^{r}+\Delta{\rm PA}^{\rm det}\,.$ (3)
Here $\Delta{\rm PA}^{a}$ is the ALDM signal defined in Eq. (2). $\Delta{\rm PA}^{\rm ion}$ is the Faraday rotation induced by the Earth’s ionosphere, which varies on timescales of hours to years. To account for unmodeled noise after ionospheric subtraction, we incorporate additional random and deterministic noise components, following common practices in pulsar timing analysis. Specifically, in addition to radiometer noise, we include extra parameters in the white noise $\Delta{\rm PA}^{w}$ to capture non-Gaussianity of observational noise, random pulse-to-pulse variation of polarization profile and imperfectly modeled ionospheric effects on short timescales. $\Delta{\rm PA}^{r}$ denotes red or temporally correlated noise, modeling variations in ISM Faraday rotation and ionospheric effects over years with a frequency-domain power spectrum. $\Delta{\rm PA}^{\rm det}$ represents long-term deterministic effects, such as ISM property changes over decades, encoded as a polynomial up to the second order (see Supplemental Material (SM) at Sec. B for details).
The ALDM signal nature can be encoded in its covariance matrix $\boldsymbol{C}^{a}$, as a two-point correlation function of $\Delta{\rm PA}^{a}$ Liu et al. (2023). In the data, $\Delta{\rm PA}^{a}$ is constructed as a vector $\{\Delta{\rm PA}_{p,n}^{a}\}$, where $p$ runs over all PPA pulsars and $n$ runs over all epochs for each of them. Inheriting from the ALDM properties Foster et al. (2021), this vector follows a multivariate Gaussian distribution with zero mean. By integrating the velocity of ALDM particles with a usual isotropic distribution peaking at the halo virial velocity, i.e. $v_{0}\sim\mathcal{O}(100)\,$km/s, we find $C_{p,n;q,m}^{a}\approx S^{2}_{a}\,\hat{C}_{p,n;q,m}^{a}\,$ Liu et al. (2023), where $S_{a}\equiv g_{a\gamma\gamma}\sqrt{\rho_{0}}/m_{a}$ is characteristic magnitude of the ALDM-induced PA residual, and
@@
reflects quadratic interplay of its Earth and pulsar terms (see Eq. (2)). $\hat{C}_{p,n;q,m}^{a}$ is essentially a correlation measure. In its expression, $t_{p,n}$ and $t^{\prime}_{p,n}=t_{p,n}-D_{p}$ denote the light reception and emission time for the $n$-th epoch of the $p$-th pulsar, with $D_{p}$ being the pulsar distance to the Earth, and their cosine functions denote temporal correlations of pulsar signals. The sinc functions ${\rm sinc}(y_{ij})\equiv\sin(y_{ij})/y_{ij}$ are all related to the pulsar term and manifest signal spatial correlations. Here, $y_{ij}\equiv|\mathbf{x}_{i}-\mathbf{x}_{j}|/l_{c}$ is a dimensionless distance parameter, and $l_{c}\equiv 1/(m_{a}v_{0})$ denotes the ALDM coherence length. These spatial correlations get suppressed when $|\mathbf{x}_{i}-\mathbf{x}_{j}|$ is significantly greater than $l_{c}$.
These correlations origin from the spacetime dependence of the ALDM field phase factor (see Eq. (1)) or equivalently the ALDM wave nature. All terms in $\hat{C}_{p,n;q,m}^{a}$ can contribute significantly to recognizing the ALDM signals Liu et al. (2023). Due to its nonrelativistic nature, the ALDM coherence length is approximately three orders of magnitude longer than $\sim 1/m_{a}$. The nanoHz ($\sim 10^{-23}\,$eV) ALDM thus has a coherence length of thousands of parsec, yielding $y_{ij}\lesssim 1$ for many PPPA pulsars. This outcome contrasts with the PTA response to the nanoHz SGWB, where the H-D curve arises from the interplay between the Earth terms, and the pulsar terms are less important since the coherence length of GWs, which is simply the inverse of their frequency, is much shorter than general pulsar distances for the current PTAs.
For the PPA analysis, the matrix $\boldsymbol{C}^{a}$ defines a signal model with full-correlation, with its diagonal ($p=q$) and off-diagonal blocks ($p\neq q$) manifesting the pulsar auto-correlation and cross-correlation, respectively. Below, we will focus on the full-correlation signal model, and as a reference, consider also the auto-correlation-only model where $\hat{C}^{a,{\rm auto}}_{p,n;q,m}=\hat{C}^{a}_{p,n;q,m}\delta_{pq}$.
@@
Here $\widetilde{\Delta{\rm PA}}\equiv\Delta{\rm PA}^{\rm obs}-\Delta{\rm PA}^{\rm ion}-\Delta{\rm PA}^{\rm det}$ is the time series of PA residuals with $\Delta{\rm PA}^{\rm ion}$ and $\Delta{\rm PA}^{\rm det}$ subtracted, where $\Delta{\rm PA}^{\rm ion}$ is derived with the package ionFR. $\boldsymbol{C}=C^{w}+C^{r}+S_{a}^{2}\hat{C}^{a}$ is the full covariance matrix that includes the contributions of the white and red noises and the ALDM signals. As the observation time span $T_{\rm obs}$ is much shorter than the ALDM coherence time ($\sim 1/(m_{a}v_{0}^{2})$), the stochastic parameters $\alpha_{\bf v}$ and $\phi_{\bf v}$ in the ALDM field have been marginalized Centers et al. (2021); Nakatsuka et al. (2023). Then by integrating out the $\Delta{\rm PA}^{\rm det}$ parameters we obtain a likelihood parameterized by $\{\theta_{n},D_{p}^{*},m_{a},S_{a}\}$, where $\theta_{n}$ denotes the set of random noise parameters, and $D_{p}^{*}$ is a pulsar distance normalized by the measured central value. To determine the spatial correlations in Eq. (4), we account for the $D_{p}^{*}$ uncertainties using priors determined by the measurement methods. For any given ALDM mass, we then perform a Bayesian analysis to constrain $S_{a}$ with this marginalized likelihood. Eventually, we convert the posterior distributions of $S_{a}$ to the limits on the ALDM Chern-Simons coupling for a given benchmark value of $\rho_{0}$. Here, we use the averaged ALDM energy density $\rho_{0}=0.4\,\textrm{GeV}/\textrm{cm}^{3}$ as the benchmark value. In addition to imposing limits, we use two Bayes factors ${\rm BF}^{\rm Full}_{\rm Null}$ and ${\rm BF}^{\rm Full}_{\rm Auto}$ to compare the full-correlation signal model with the null-signal model and to differentiate between different signal models. The latter is particularly crucial for understanding the nature of the derived limits (see SM at Secs. A and B for details).
Figure 1: 95% upper limits on the ALDM characteristic signal strength $S_{a}$ as a function of its mass $m_{a}$, for six representative PPPA MSPs. In each panel, the black solid line represents the single-pulsar limits with full noise modeling in Eq. (3), while the colored lines show the limits excluding individual noise components: additional white noise (blue), red noise (orange), linear and quadratic deterministic noise (purple), and ionospheric effects (green). The difference between the black and colored lines reflects the contribution of each noise component to the analysis. As a reference, we present observation period $T_{\rm obs}$ of each pulsar, and time scales of solar cycle (11 years), 1 year and 1 day, as colored and grey vertical dotted lines, respectively. The timescale and $m_{a}$ are related by $m_{a}=2\pi/T$.
@@ IV PPA results and discussions
The PPTA DR3 Zic et al. (2023) includes the measurements of Stokes parameters for 32 MSPs in the radio band, with a period from 2004 to 2022. We reprocess all data taken with the Ultra-Wideband Low receiver system, with the most up-to-date polarization calibration model to improve the measurement quality Hobbs et al. (2020), and correct for static Faraday rotations caused by the ISM using the rotation measures provided by the Australia Telescope National Facility pulsar catalogue Manchester et al. (2005). We then select 22 MSPs for the PPPA construction whose signal-to-noise ratio of linear polarization intensity is relatively high in the considered frequency band ($2.6-3.6\,$GHz). For these MSPs, we align the observation in each epoch with the PPTA template of total intensity Zic et al. (2023); Dai et al. (2015). We then derive the standard PA profiles, subtract them from the data, and perform a weighted average over the phase bins, to generate PA residual time series for unbiased signal estimation. (see SM at Sec. A for details).
@@
We then test the noise model for individual MSPs using the likelihood in Eq. (5), with the ALDM signal turned off. Compared to a minimal model including only the observation error of $\Delta{\rm PA}^{\rm obs}$ and a constant deterministic residual Liu et al. (2023), the full noise model developed in this study shows a remarkable improvement in fitting the data for the majority of the 22 PPPA MSPs. The most significant improvements are observed for J0437-4715, J1713+0747 and J1909-3744. The contribution of each new component vary greatly among different pulsars. The analysis suggests that the imperfect modeling of ionospheric Faraday rotation may account for the additional white noise and even red noise observed for some specific pulsars, while slow change of the ISM properties over decades could explain the observed deterministic linear and quadratic variations. This outcome fully demonstrates the effectiveness of incorporating new noise components in modeling the observed PA residuals in Eq. (3).
@@
We demonstrate in Fig. 1 the 95% upper limits on the ALDM characteristic signal strength $S_{a}$ as a function of $m_{a}$, for six representative PPPA MSPs individually. For the mass range of $m_{a}\gtrsim 1/T_{\rm obs}$, the $S_{a}$ limits are relatively flat. A best limit as strong as 0.10.10.1\,0.1deg or even stronger can be achieved for many of these pulsars including J0437-4715, J1744-1134, J1022+1001, etc. As the ALDM mass decreases to below $1/T_{\rm obs}$, the limits deteriorate approximately with $m_{a}^{-3}$, consistent with the expectation from integrating $\Delta{\rm PA}^{\rm det}$ as a second-order polynomial expansion Hazboun et al. (2019). The ionospheric subtraction effectively removes sharp peaks around $\sim 1\,$yr or/and low-frequency excess for J0437-4715, J1600-3053 and J1744-1134. The incorporation of additional white noise component enhances the limits for six pulsars uniformly, and of red noise component improves intermediate- and low-frequency limits for J0437-4715 and J1909-3744 notably. These analyses show that, due to the variance of noise nature over pulsars, the advanced noise model in Eq. (3) is essential for getting robust limits on the ALDM signal strength (see SM at Sec. B for details).
@@
Figure 2: 95% upper limits of the PPPA on the ALDM Chern-Simons coupling $g_{a\gamma\gamma}$ as a function of its mass $m_{a}$, with $\rho_{0}=0.4\,\textrm{GeV}/\textrm{cm}^{3}$. The blue and red curves represent the limits obtained with the auto-correlation-only and full-correlation signal models, respectively, with (solid) and without (dashed) the ionospheric subtraction. As a reference, the smoothened constraints from the observation of PA variation in Crab Nebula with POLARBEAR Adachi et al. (2024), the analysis of PPTA data of DR1 and Crab Nebula data from QUIJOTE Castillo et al. (2022), the measurement of the CMB polarization with SPT-3G Ferguson et al. (2022), and the investigation of PLANCK mission on the ALDM-induced washout effects during recombination Fedderke et al. (2019) are presented as cyan, yellow, green, and purple dashed lines. The limits from the CAST experiment Anastassopoulos et al. (2017) and the observations of SN1987A Payez et al. (2015) and X-ray spectral distortion in quasar H1821+643 Reynés et al. (2021) are also presented as black dashed lines. The shaded regions above these lines indicate the range of $m_{a}$ and $g_{a\gamma\gamma}$ that they have excluded. The vertical dotted lines denote the reference time scales, including the PPPA observation period $T_{\rm obs}\approx 18~{}{\rm yrs}$, solar cycle, one year and one day. In the lower sub-panel, the Bayes factors are displayed in two scenarios: the full-correlation signal model against the null-signal model (red) and the full-correlation signal model against the auto-correlation-only signal model (black).
@@
The final PPPA limits on the ALDM Chern-Simons coupling $g_{a\gamma\gamma}$ are presented in Fig. 2 as a function of its mass $m_{a}$, for both full-correlation and auto-correlation-only signal models. These limits exhibit a linear dependence on $m_{a}$ for $m_{a}\gtrsim 1/T_{\rm obs}$, while scale as $m_{a}^{-2}$ for $m_{a}\lesssim 1/T_{\rm obs}$. This feature should be expected from the scaling behavior of the $S_{a}$ limits depicted in Fig. 1. Moreover, one can see that the ionospheric subtraction softens the peak at the time scale of 1 year in these limits. These limits, however, start to oscillate strongly as the time scale decreases to one day or shorter, resulting in some sharp peaks for $m_{a}\gtrsim 10^{-19.4}\,$eV.

### v2

@@ I Introduction
Pulsars are also known as excellent astronomical linear polarization sources. Their light can be linearly polarized to wellabove $10\%$ of total intensity. The PTA programs usually measure the pulsar polarization profiles for calibrating the relevant observations Osłowski et al. (2013). Vast amounts of polarization data thus exist or are expected to be collected in the active PTAs and other pulsar-monitoring programs. This has motivated the suggestion of developing pulsar polarization arrays (PPAs) Liu et al. (2023), to explore astrophysics and fundamental physics with a correlated polarization signal across the pulsars. One representative scientific case is detecting ultralight axion-like dark matter (ALDM).
@@
In this letter, we conduct the first-ever PPA analysis to search for the ultralight ALDM, using the polarization data of 22 MSPs from PPTA Data Release 3 (DR3) Zic et al. (2023); Reardon et al. (2023). We denote this PPA as the Parkes PPA (PPPA) for the convenience of discussions. To accomplish this task, we develop an advanced analysis framework tailored specifically for constructing real pulsar PA residual time series, modeling instrumental and astronomical noises and searching for signal correlation across the arrayed pulsars. Using this Bayesian framework, we demonstrate that the PPPA provides the most stringent constraints on the ALDM Chern-Simons coupling within the relevant mass range. Notably, we demonstrate for the first time that the pulsar cross-correlation plays a crucial role in investigating the nature of the derived limits.
@@ II ALDM induced cosmic birefringence
$\times\cos[m_{a}(t-\mathbf{v}\cdot\mathbf{x})+\phi_{\mathbf{v}}]\,$,
where $\rho(\mathbf{x})$ is ALDM energy density at position $\mathbf{x}$, $m_{a}$ is ALDM mass, $\mathbf{v}$ denotes velocity in a latticed phase space $\Omega$, and $\Delta v$ is lattice spacing. $f_{\mathbf{x}}(\mathbf{v})$ represents DM velocity distribution at $\mathbf{x}$, which is assumed to be spatially uniform and isotropic, with the speed $|\mathbf{v}|$ peaking at the halo’s virial velocity, i.e. $v_{0}\sim\mathcal{O}(100)\,$km/s. The profile of $\rho(\mathbf{x})$ is determined by Schrödinger-Poisson equations. As shown in Schive et al. (2014), the halo self-gravitation leads to the formation of a central solitonic core within galaxies. Since most PPPA MSPs are only $\lesssim\mathcal{O}(10^{3})$ parsec away from the Earth, we simply assume $\rho(\mathbf{x})$ to be uniform, i.e., $\rho(\mathbf{x})=\rho_{0}$ foo , with its value to be determined by galactic halo measurements. The parameters $\alpha_{\mathbf{v}}\in(0,+\infty)$ and $\phi_{\mathbf{v}}\in(0,2\pi)$ represent random amplitude and phase, respectively, following Rayleigh and uniform distributions Foster et al. (2018). For a specific realization of $\{\alpha_{\mathbf{v}},\phi_{\mathbf{v}}\}$, the ALDM field $a(\mathbf{x},t)$ exhibits stochastic spacetime dependence Centers et al. (2021), where coherence gets lost for a temporal interval $\gg\tau_{c}$ or a spatial interval $\gg l_{c}$, with $\tau_{c}\sim 1/(m_{a}v_{0}^{2})$ and $l_{c}\sim 1/(m_{a}v_{0})$ denoting the coherent time and coherent length (i.e. de Broglie wavelength) of this field. For the mass range of ultralight ALDM, the observation period $T_{\rm obs}$ for existing PTA programs is significantly shorter than $\tau_{c}$, leading to coherent temporal evolution of $\cos(m_{a}t+\ldots)$ within this period as shown in Eq. (1). However, the distances of pulsars, whether to another pulsar or to the Earth, may exceed $l_{c}$, resulting in stochastic variation of the ALDM field over such distances.
@@
where the two terms are dubbed “pulsar” and “Earth” terms. The $\Delta{\rm PA}^{a}$ magnitude can be characterized by a parameter $S_{a}\equiv g_{a\gamma\gamma}\sqrt{\rho_{0}}/m_{a}$ then. For $m_{a}\sim 10^{-22}\,$eV and $\rho_{0}\sim 0.1\,\textrm{GeV}/\textrm{cm}^{3}$, this implies that a signal of $\Delta{\rm PA}^{a}\sim 0.1\,$deg needs the Chern-Simons coupling to be $g_{a\gamma\gamma}\sim 10^{-13}\,\textrm{GeV}^{-1}$. Note, while the SGWB-induced timing residual in the PTA is also determined by so-called “pulsar” and “Earth” terms, unlike the CB effect, it arises from a cancellation between contraction and expansion of the light path as the GWs pass by. Moreover, the CB is distinguished from the Faraday rotation in interstellar medium (ISM) and Earth’s ionosphere as the latter effect is frequency dependent and relies on the length of light path aligned with the direction of magnetic field.
@@
In last decade, a few attempts have been made to explore temporal variations of pulsar polarization Weisberg et al. (2010); Yan et al. (2011); Wahl et al. (2022). The PA residuals of individual pulsars have been suggested for testing the ALDM theory Liu et al. (2020); Caputo et al. (2019). However, a systematic analysis framework of the PPA, which emphasizes the data correlation across the pulsars, remains absent. To address this, we model the PA residuals as:
$\Delta{\rm PA}^{\rm obs}=\Delta{\rm PA}^{a}+\Delta{\rm PA}^{\rm ion}+\Delta{\rm PA}^{w}+\Delta{\rm PA}^{r}+\Delta{\rm PA}^{\rm pol}\,.$ (3)
Besides the stochastic signal $\Delta{\rm PA}^{a}$ defined in Eq. (2), there are additional components categorized as deterministic and random. The deterministic components include the Faraday rotation induced by the Earth’s ionosphere, $\Delta{\rm PA}^{\rm ion}$, which varies over timescales from hours to years and is derived with the package ionFR. The random contributions consist of white noise, $\Delta{\rm PA}^{w}$, which incorporates radiometer noise and accounts for random pulse-to-pulse variations in polarization profiles, along with unmodeled ionospheric effects that persist after the ionFR subtraction on shorter timescales. Red noise, $\Delta{\rm PA}^{r}$, is also random and captures variations in ISM Faraday rotation or fluctuations in dispersion measure and rotation measure and unmodeled ionospheric effects over years using a frequency-domain power spectrum. Additionally, we model long-term effects observed in the data with a polynomial of up to second order, denoted by $\Delta{\rm PA}^{\rm pol}$, which may account for slow stochastic changes in ISM properties over decades. (see Supplemental Material (SM) at Sec. B for details).
The ALDM signal nature can be encoded in its covariance matrix $\boldsymbol{C}^{a}$, as a two-point correlation function of $\Delta{\rm PA}^{a}$ Liu et al. (2023). In the data, $\Delta{\rm PA}^{a}$ is constructed as a vector $\{\Delta{\rm PA}_{p,n}^{a}\}$, where $p$ runs over all PPA pulsars and $n$ runs over all epochs for each of them. Inheriting from the ALDM properties Foster et al. (2021), this vector follows a multivariate Gaussian distribution with zero mean. By integrating the velocity of ALDM particles with a delta-function approximation for the speed at $|\mathbf{v}|=v_{0}$, we find $C_{p,n;q,m}^{a}\approx S^{2}_{a}\,\hat{C}_{p,n;q,m}^{a}\,$ Liu et al. (2023),where
@@
reflects quadratic interplay of its Earth and pulsar terms (see Eq. (2)). $\hat{C}_{p,n;q,m}^{a}$ is essentially a correlation measure. In its expression, $t_{p,n}$ and $t^{\prime}_{p,n}=t_{p,n}-D_{p}$ denote the light reception and emission time for the $n$-th epoch of the $p$-th pulsar, with $D_{p}$ being the pulsar distance to the Earth, and their cosine functions denote temporal correlations of pulsar signals. The sinc functions ${\rm sinc}(y_{ij})\equiv\sin(y_{ij})/y_{ij}$ are all related to the pulsar term and manifest signal spatial correlations. Here, $y_{ij}\equiv|\mathbf{x}_{i}-\mathbf{x}_{j}|/l_{c}$ is a dimensionless distance parameter. These spatial correlations get suppressed when $|\mathbf{x}_{i}-\mathbf{x}_{j}|$ is significantly greater than the coherence length $l_{c}$.
These correlations originate from the spacetime dependence of the ALDM field phase factor (see Eq. (1)) or equivalently the ALDM wave nature. All terms in $\hat{C}_{p,n;q,m}^{a}$ can contribute significantly to recognizing the ALDM signals Liu et al. (2023). Due to its nonrelativistic nature, the ALDM de Broglie wavelength is approximately three orders of magnitude longer than its Compton wavelength ($\sim 1/m_{a}$). The nanoHz ($\sim 10^{-23.4}\,$eV) ALDM thus has a de Broglie wavelength $\sim 1.5\,$kpc, yielding $y_{ij}\lesssim 1$ for many PPPA pulsars. This outcome contrasts with the PTA response to the nanoHz SGWB, where the H-D curve arises from the interplay between the Earth terms, and the pulsar terms are less important since the de Broglie wavelength of GWs is simply their Compton wavelength and hence much shorter than general pulsar distances for the current PTAs.
For the PPA analysis, the matrix $\boldsymbol{C}^{a}$ defines a signal model with full correlation, with its diagonal ($p=q$) and off-diagonal blocks ($p\neq q$) manifesting the pulsar autocorrelation and cross-correlation, respectively. Below, we will focus on the full-correlation signal model, and as a reference, consider also the autocorrelation-only model where $\hat{C}^{a,{\rm auto}}_{p,n;q,m}=\hat{C}^{a}_{p,n;q,m}\delta_{pq}$.
@@
Here $\widetilde{\Delta{\rm PA}}\equiv\Delta{\rm PA}^{\rm obs}-\Delta{\rm PA}^{\rm ion}-\Delta{\rm PA}^{\rm pol}$ is the time series of PA residuals with $\Delta{\rm PA}^{\rm ion}$ and $\Delta{\rm PA}^{\rm pol}$ subtracted. $\boldsymbol{C}=C^{w}+C^{r}+S_{a}^{2}\hat{C}^{a}$ is the full covariance matrix that includes the contributions of the white and red noises and the ALDM signals. For the parameter $S_{a}$, we set the average ALDM energy density $\rho_{0}=0.4\,\textrm{GeV}/\textrm{cm}^{3}$ Salucci et al. (2010) as the benchmark, since the Milky Way inner region ($\sim 20\,$kpc), where its rotation curve is measured for determining the local DM density Nesti and Salucci (2013), is considerably larger than $l_{c}$ for the mass range considered below. Then by integrating out the $\Delta{\rm PA}^{\rm pol}$ parameters we obtain a likelihood parameterized by $\{\theta_{n},D_{p}^{*},m_{a},S_{a}\}$, where $\theta_{n}$ denotes the set of random noise parameters, and $D_{p}^{*}$ is a pulsar distance normalized by the measured central value. To determine the spatial correlations in Eq. (4), we account for the $D_{p}^{*}$ uncertainties using priors determined by the measurement methods. For any given ALDM mass, we then perform a Bayesian analysis to constrain $S_{a}$ with this marginalized likelihood. Eventually, we convert the posterior distributions of $S_{a}$ to the limits on the ALDM Chern-Simons coupling for the given benchmark value of $\rho_{0}$. In addition to imposing limits, we use two Bayes factors ${\rm BF}^{\rm Full}_{\rm Null}$ and ${\rm BF}^{\rm Full}_{\rm Auto}$ to compare the full-correlation signal model with the null-signal model and to differentiate it from the autocorrelation-only signal model, respectively. The latter is particularly crucial for understanding the nature of the derived limits and recognizing the real signals in the data (see SM at Secs. A and B for details).
Figure 1: 95% upper limits on the ALDM characteristic signal strength $S_{a}$ as a function of its mass $m_{a}$, for six representative PPPA MSPs. In each panel, the black solid line represents the single-pulsar limits with full noise modeling in Eq. (3), while the colored lines show the limits excluding individual noise components: additional white noise (blue), red noise (orange), linear and quadratic deterministic noise (purple), and ionospheric effects (green). The difference between the black and colored lines reflects the contribution of each noise component to the analysis. As a reference, we present $T_{\rm obs}$ of each pulsar, and timescales of solar cycle (11 years), one year and one day, as colored and gray vertical dotted lines, respectively. The timescale and $m_{a}$ are related by $m_{a}=2\pi/T$.
@@ IV PPA results and discussions
The PPTA DR3 Zic et al. (2023) includes the measurements of Stokes parameters for 32 MSPs in the radio band, with a period from 2004 to 2022. We reprocess all data taken with the Ultra-Wideband Low receiver system, with the most up-to-date polarization calibration model to improve the measurement quality Hobbs et al. (2020), and correct for static Faraday rotations caused by the ISM using the rotation measures provided by the Australia Telescope National Facility pulsar catalog Manchester et al. (2005). We then select 22 MSPs for the PPPA construction whose signal-to-noise ratio of linear polarization intensity is relatively high in the considered frequency band (2.63.62.6-3.6\,GHz). For these MSPs, we align the observation in each epoch with the PPTA template of total intensity Zic et al. (2023); Dai et al. (2015). We then derive the standard PA profiles, subtract them from the data, and perform a weighted average over the phase bins, to generate PA residual time series for unbiased signal estimation (see SM at Sec. A for details).
@@
We then test the noise model for individual MSPs using the likelihood in Eq. (5), with the ALDM signal turned off. To achieve this, we introduce a set of Bayes factors as a metric to assess the improvements in noise modeling by adding additional noise components one at a time to the minimal noise model, ultimately leading to the full noise model (see Table II at Sec. B of SM for details). Compared to the minimal model including only the observation error of $\Delta{\rm PA}^{\rm obs}$ and a constant deterministic residual Liu et al. (2023), the full noise model shows a remarkable improvement in fitting the data for the majority of the 22 PPPA MSPs. The most significant improvements are observed for J0437-4715, J1713+0747 and J1909-3744. The contribution of each new component vary greatly among different pulsars. The analysis suggests that the imperfect modeling of ionospheric Faraday rotation may account for the additional white noise and even red noise observed for some specific pulsars, while slow change of the ISM properties over decades could explain the observed deterministic linear and quadratic variations. This outcome fully demonstrates the effectiveness of incorporating new noise components in modeling the observed PA residuals in Eq. (3).
@@
We demonstrate in Fig. 1 the 95% upper limits on the ALDM characteristic signal strength $S_{a}$ as a function of $m_{a}$, for six representative PPPA MSPs individually. For the mass range of $m_{a}\gtrsim 1/T_{\rm obs}$, the $S_{a}$ limits are relatively flat. A best limit as strong as 0.10.1\,deg or even stronger can be achieved for many of these pulsars including J0437-4715, J1744-1134, J1022+1001, etc. As the ALDM mass decreases to below $1/T_{\rm obs}$, the limits deteriorate approximately with $m_{a}^{-3}$, consistent with the expectation from integrating $\Delta{\rm PA}^{\rm pol}$ as a second-order polynomial expansion Hazboun et al. (2019). The ionospheric subtraction effectively removes sharp peaks around $\sim 1\,$yr or/and low-frequency excess for J0437-4715, J1600-3053 and J1744-1134. The incorporation of additional white noise component enhances the limits for six pulsars uniformly, and of red noise component improves intermediate- and low-frequency limits for J0437-4715 and J1909-3744 notably. These analyses show that, due to the variance of noise nature over pulsars, the advanced noise model in Eq. (3) is essential for getting robust limits on the ALDM signal strength (see SM at Sec. B for details).
@@
Figure 2: 95% upper limits of the PPPA on the ALDM Chern-Simons coupling $g_{a\gamma\gamma}$ as a function of its mass $m_{a}$, with $\rho_{0}=0.4\,\textrm{GeV}/\textrm{cm}^{3}$. The blue and red curves represent the limits obtained with the autocorrelation-only and full-correlation signal models, respectively, with (solid) and without (dashed) the ionospheric subtraction. As a reference, the smoothened constraints from the observation of PA variation in Crab Nebula with POLARBEAR Adachi et al. (2024), the analysis of PPTA data of DR1 and Crab Nebula data from QUIJOTE Castillo et al. (2022), the measurement of the CMB polarization with SPT-3G Ferguson et al. (2022), and the investigation of PLANCK mission on the ALDM-induced washout effects during recombination Fedderke et al. (2019) are presented as cyan, yellow, green, and purple dashed lines. The limits from the CAST experiment Anastassopoulos et al. (2017) and the observations of SN1987A Payez et al. (2015) and x-ray spectral distortion in quasar H1821+643 Reynés et al. (2021) are also presented as black dashed lines. The shaded regions above these lines indicate the range of $m_{a}$ and $g_{a\gamma\gamma}$ that they have excluded. The vertical dotted lines denote the reference timescales, including the PPPA observation period $T_{\rm obs}\approx 18~{\rm yrs}$, solar cycle, one year and one day. In the lower subpanel, the Bayes factors are displayed in two scenarios: the full-correlation signal model against the null-signal model (red) and the full-correlation signal model against the autocorrelation-only signal model (black).
@@
The final PPPA limits on the ALDM Chern-Simons coupling $g_{a\gamma\gamma}$ are presented in Fig. 2 as a function of its mass $m_{a}$, for both full-correlation and autocorrelation-only signal models. These limits exhibit a linear dependence on $m_{a}$ for $m_{a}\gtrsim 1/T_{\rm obs}$, while scale as $m_{a}^{-2}$ for $m_{a}\lesssim 1/T_{\rm obs}$. This feature should be expected from the scaling behavior of the $S_{a}$ limits depicted in Fig. 1. Moreover, one can see that the ionospheric subtraction softens the peak at the timescale of 1 year in these limits. These limits, however, start to oscillate strongly as the timescale decreases to one day or shorter, resulting in some sharp peaks for $m_{a}\gtrsim 10^{-19.4}\,$eV.